# HoTPP Benchmark: Are We Good at the Long Horizon Events Forecasting?

### v1

@@ C.2 Metrics
MTPP models are typically evaluated by the quality of the next event prediction. These metrics usually include mean absolute error (MAE) or root mean square error (RMSE) for time shifts and accuracy or error rate for label quality evaluation. Some works also evaluate test set likelihood as predicted by the model. We do not measure this quantity, as it is intractable for MAE-CE and Next-K models. Previous works have made progress in evaluating long-horizon prediction [mei2019imputing, xue2023easytpp] with optimal transport distance (OTD). We evaluate these metrics, as well as the novel T-mAP metric, which fixes the flaws of previous metrics, as discussed in Section 3. We set the maximum time error $\delta$ for the T-mAP metric twice the cost of the OTD because when the model predicts an incorrect label, OTD removes the prediction and adds the ground truth event with the total cost equal to $2C$.
@@ C.3 Backbones
We use two types of architectures in our experiments. The first one is the GRU network[cho2014gru], which is one of the top-performing neural architectures in the event sequences domain [babaev2022coles]. We also implement a continuous-time LSTM (CT-LSTM) from the NHP method [mei2017nhp]. As CT-LSTM requires a special loss function and increases training time, we apply it only in the NHP method. In other cases, we prefer GRU. Another advantage of GRU is that its output is equal to its hidden state, which simplifies intermediate hidden state estimation for autoregressive inference starting from the middle of a sequence.
@@ C.4 Datasets
For the first time, we combine domains of financial transactions, social networks, and medical records in a single evaluation benchmark. Specifically, we provide evaluation results on a transactional dataset (AgePred) [babaev2022coles], MIMIC-IV [johnson2020mimic], social networks datasets (Retweet [zhao2015seismic] and StackOverflow [jure2014snap]), and Amazon dataset from the area of recommender systems [jure2014amazon]. These datasets have different underlying processes. For example, social network datasets are subject to cascades [zhao2015seismic]. Medical records show repetitive patterns and transactional data resembles daily activities with some regularity and large uncertainty.
The dataset statistics are presented in Table 2. Transactions have the largest average sequence length and the largest number of classes. MIMIC-IV has the largest number of sequences. Retweet is medium size and Amazon with StackOverflow can be considered small datasets.
@@
MIMIC-IV is a publicly available electronic health record database, that includes patient diagnoses, lab measurements, procedures, and treatments. We use data preprocessing from EventStreamGPT to construct intermediate representations in EFGPT format. The following entities are obtained: subjects data frame with time-independent patient records, events data frame with lists of event types happening to subjects at specific timestamps, and measurements data frame containing specific values of time-dependent measurements and linked to events data frame.
We join events and measurements and form sequences of events for subjects. The classification labels can be generally separated into diagnoses in the form of ICD codes and types of events, such as admissions, procedures, measurements, etc. The main issue is the inability to use either diagnoses only or event types only. In the former case, the sequences become too short, while in the latter, the sequences become too repetitive, with periodic events, such as treatment start/finish, dominating the data.
However, using both diagnoses and event types simultaneously presents another problem, as diagnosis designations are distributed sparsely in the constant stream of repeat procedures. This imbalanced class distribution causes poor performance. Therefore, we further filter the data by removing duplicate events between diagnoses. This way, we can retain useful treatment data and preserve class balance.
The final labels are formed by converting ICD-9 and ICD-10 codes to ICD-10 chapters using General Equivalence Mapping (GEM) and adding event types as additional classes. The conversion is needed because the number of ICD codes is too large to use them as classes directly. We also address the issue of multiple diagnoses being designated in a single event by sorting timestamps for reproducibility.
@@ Appendix D Training details
We trained each model with the Adam optimizer, learning rate set to $0.001$, and learning rate scheduler, decreasing learning rate by 20
We used NVIDIA V100 and A100 GPUs for computation. Each method was trained with a single GPU. The training time depends on the dataset and method. It varies from 5 minutes for RMTPP on StackOverflow to 40 minutes for the same method on Transactions or even 15 hours for the NHP method on the Transactions dataset. Multiseed evaluation takes 5 times more time to finish.
@@ Appendix E Hyperparameters
We analyzed the behavior of OTD and T-mAP metrics depending on their parameters. The results are shown in Figure 9. OTD cost has little impact on methods ranking but largely affects the metric scale. The choice of $\delta$ parameter in T-mAP can, in some cases, affect methods ranking, but if it is not too small and too large, the metric demonstrates stable ranking. We set the $\delta$ parameter to about 10-30% of the horizon duration.
Evaluation hyperparameters are presented in Table 3. Dataset-specific training parameters are presented in Table 4.
@@
Dataset Maximum Length OTD prefix size OTD Cost $C$ T-mAP horizon T-mAP $\delta$ Transactions 16 5 1 7 (week) 2 MIMIC-IV 16 5 2 28 (days) 4 Retweet 16 10 15 180 (seconds) 30 Amazon 16 5 1 10 (unk. unit) 2 StackOverflow 16 10 1 10 (minutes) 2
@@ Appendix F Labels Entropy Degradation
We omitted some datasets in the main part of the paper for simplicity. In Fig. 7 one can see the dependency of the labels distribution entropy depending on step size for all datasets.
Figure 7: Labels distribution entropy depending on the generated sequence position.
@@ Appendix G The Optimal Sequence Length
We omitted some datasets in the main part of the paper for simplicity. In Fig. 8 you can see the dependency of the labels distribution entropy depending on step size for all datasets.
@@
Figure 9: The dependency of metric values on the metric parameter.

### v2

@@ C.2 Metrics
MTPP models are typically evaluated based on the accuracy of next-event predictions. Common metrics include mean absolute error (MAE) or root mean square error (RMSE) for time shifts, and accuracy or error rate for label predictions. Some studies also assess test set likelihood as predicted by the model, but we do not include this measure as it is intractable for IFTPP and Next-K models. Previous works have advanced long-horizon evaluation using optimal transport distance (OTD) (Mei et al., 2019; Xue et al., 2023). In addition to these metrics, we evaluate the novel T-mAP metric, which addresses the shortcomings of previous metrics, as discussed in Section 3.
@@ C.3 Backbones
We use three types of architectures in our experiments. The first is the GRU network (Cho et al., 2014), which is one of the top-performing neural architectures for event sequences (Babaev et al., 2022). We also implement the continuous-time LSTM (CT-LSTM) from the NHP method (Mei & Eisner, 2017). Since CT-LSTM requires a specialized loss function and increases training time, we use it exclusively with the NHP method, preferring GRU in other cases. An additional advantage of GRU is that its output is equal to its hidden state, simplifying the estimation of intermediate hidden states for autoregressive inference starting from the middle of a sequence. Lastly, we implement the AttNHP continuous-time transformer model (Yang et al., 2022).
@@ C.4 Datasets
For the first time, we combine domains such as financial transactions, social networks, and medical records into a single evaluation benchmark. Specifically, we provide evaluation results on a transactional dataset (Babaev et al., 2022), the MIMIC-IV medical dataset (Johnson et al., 2023), and social network datasets (Retweet (Zhao et al., 2015), StackOverflow (Jure, 2014), and Amazon (Jianmo, 2018)). These datasets represent diverse underlying processes: social network data is influenced by cascades (Zhao et al., 2015), medical records exhibit repetitive patterns, and transactional data reflects daily activities, combining regularity with significant uncertainty.
The dataset statistics are presented in Table 2. The Transactions dataset has the longest average sequence length and the largest number of classes, while MIMIC-IV contains the highest number of sequences. Retweet is of medium size, whereas Amazon and StackOverflow can be considered small datasets.
@@
MIMIC-IV is a publicly available electronic health record database that includes patient diagnoses, lab measurements, procedures, and treatments. We leverage the data preprocessing pipeline from EventStreamGPT to construct intermediate representations in EFGPT format. This process yields three key entities: a subjects data frame containing time-independent patient records, an events data frame listing event types occurring to subjects at specific timestamps, and a measurements data frame with time-dependent measurement values linked to the events data frame.
We combine events and measurements to create sequences of events for each subject. The classification labels are generally divided into two categories: diagnoses, represented by ICD codes, and event types, such as admissions, procedures, and measurements. A key challenge is that relying solely on diagnoses results in sequences that are too short, while using only event types leads to highly repetitive sequences dominated by periodic events, such as treatment start and finish.
However, using both diagnoses and event types together introduces another issue: diagnoses are sparsely distributed within a constant stream of repeated procedures, leading to imbalanced class distributions and poor performance. To mitigate this, we filter the data by removing duplicate events between diagnoses, allowing us to retain useful treatment data while preserving class balance.
The final labels are created by converting ICD-9 and ICD-10 codes to ICD-10 chapters using General Equivalence Mapping (GEM) and adding event types as additional classes. This conversion is necessary because the number of ICD codes is too large to use them directly as classes. Additionally, we address the issue of multiple diagnoses occurring in a single event by sorting timestamps for reproducibility.
C.5 Performance improvements
The HoTPP benchmark provides highly optimized training and inference procedures for efficient evaluation on datasets containing up to tens of millions of events. First, we implement parallel RNN inference, which reuses computations when inference is initiated from multiple starting points within a sequence. Additionally, we optimize the code and apply PyTorch Script to the CT-LSTM model from NHP and the ODE model. CT-LSTM from EasyTPP and the ODE model from the original repository serve as baselines. The timing results are presented in Table 3. Experiments were conducted using a synthetic batch with a size of 64, a sequence length of 100, and an embedding dimension of 64. Evaluation was performed on a single Nvidia RTX 4060 GPU. The results show that HoTPP is 17 times faster at RNN autoregressive inference compared to simple prefix extension. HoTPP also accelerates CT-LSTM and ODE by 4 to 8 times. These optimizations significantly extend the applicability of the implemented methods to larger-scale datasets [^8] Experiments are documented in the “notebooks” folder of the HoTPP repository. Additionally, HoTPP offers a GPU implementation of the Hungarian algorithm, which also finds applications in computer vision.
Table 3: HoTPP computation speed improvements in terms of seconds per batch.
Implementation Study
RNN autoreg. inference CT-LSTM inference CT-LSTM train ODE inference ODE train
Baseline 2.44 0.0158 0.0519 0.08 0.162
HoTPP 0.14 0.0025 0.0276 0.01 0.048
@@ Appendix D Training details
We trained each model using the Adam optimizer with a learning rate of 0.001 and a scheduler that reduces the learning rate by 20% after every 5 epochs. Gradient clipping was applied, with a maximum L2-norm set to 1.
We performed computations on NVIDIA V100 and A100 GPUs, with some smaller experiments conducted on an Nvidia RTX 4060. Each method was trained on a single GPU. The training time varied depending on the dataset and method, ranging from 5 minutes for RMTPP on the StackOverflow dataset to 40 minutes for the same method on the Transactions dataset, and up to 15 hours for NHP on Transactions. Multi-seed evaluations took approximately five times longer to complete.
@@ Appendix E Hyperparameters
We analyzed the behavior of OTD and T-mAP metrics depending on their parameters. The results are shown in Figure 9. The OTD cost has little effect on the ranking of methods but significantly influenced the metric scale. The choice of the $\delta$ parameter in T-mAP can, in some cases, affect method rankings, but as long as $\delta$ is not set too small or too large, the metric demonstrates stable rankings. We set the $\delta$ parameter to approximately 10-30% of the horizon duration.
Evaluation hyperparameters are presented in Table 4. Dataset-specific training parameters are listed in Table 5.
@@
Dataset Maximum Length OTD prefix size OTD Cost $C$ T-mAP horizon T-mAP $\delta$ Transactions 32 5 1 7 (week) 2 MIMIC-IV 32 5 2 28 (days) 4 Retweet 32 10 15 180 (seconds) 30 Amazon 32 5 1 10 (unk. unit) 2 StackOverflow 32 10 1 10 (minutes) 2
@@
Figure 9: The dependency of metric values on the metric parameter.
@@ Appendix F Labels Entropy Degradation
For simplicity, some datasets were omitted in the main part of the paper. In Figure 10, we show the dependency of label distribution entropy on step size for all datasets.
Figure 10: Entropy of label distributions as a function of the position in the generated sequence.
@@ Appendix G The Optimal Sequence Length
For simplicity, some datasets were omitted from the main part of the paper. Figure 11 illustrates the relationship between label distribution entropy and step size across all datasets.